OpenStudy (anonymous):

ok this is tough... Simplify: ln((2xy^3))/(sqaure root(z))

myininaya (myininaya):

do you mean expand? \[\ln(\frac{2xy^3}{\sqrt{x}})=\ln(2xy^3)-\ln(\sqrt{x})=\ln(2)+\ln(x)+\ln(y^3)-\ln(x^\frac{1}{2})\] \[=\ln(2)+\ln(x)+3\ln(y)-\frac{1}{2}\ln(x)\]

OpenStudy (anonymous):

is that what simplify means? all it says is simplify

myininaya (myininaya):

lol i don't know what simplify means let me get satellite

myininaya (myininaya):

oops that second x was suppose to be a z

OpenStudy (anonymous):

haha yeah i caught that

myininaya (myininaya):

\[\ln(\frac{2xy^3}{\sqrt{z}})=\ln(2xy^3)-\ln(\sqrt{z})=\ln(2)+\ln(x)+\ln(y^3)-\ln(z^\frac{1}{2})\] \[=\ln(2)+\ln(x)+3\ln(y)-\frac{1}{2}\ln(z)\]

OpenStudy (anonymous):

Thank you so much!

myininaya (myininaya):

but hey i don't know what simplify means lol so i told satellite to come here so he can give his opinion

OpenStudy (anonymous):

ok gosh youre so helpful

OpenStudy (anonymous):

this one?

myininaya (myininaya):

yes we don't know what simplify means

OpenStudy (anonymous):

yes sir!

myininaya (myininaya):

and simplify is your favorite word i just thought we we were to expand because thats the only i can see to do

OpenStudy (anonymous):

ok you did it. this is not called "simplifying" however, but you have the right answer

OpenStudy (anonymous):

sometimes you are asked to write something as a single log sometimes you are asked to write something as the sum and product of logs this one is the second one

myininaya (myininaya):

i don't like the word simplify anymore

OpenStudy (anonymous):

there is absolutely nothing "simpler" about one form or the other. depends on what your aim is and frankly if your math teacher wrote "simplify" in this case, rather than saying what he or she wanted, they should not be teaching math

OpenStudy (anonymous):

haha thanks! thats my Texas A&M calc prof for ya there...

myininaya (myininaya):

lol

OpenStudy (anonymous):

ok can i ask you one more quick question?

myininaya (myininaya):

what is it

OpenStudy (anonymous):

evaluate: cos^-1 (-square root(3/2))

myininaya (myininaya):

\[\cos^{-1}(-\sqrt{\frac{3}{2}}) \] that right?

OpenStudy (anonymous):

oh no thats all right except its negative sqaure root of 3 divided by 2

OpenStudy (anonymous):

i will bet it is \[\cos^{-1}(\frac{\sqrt{3}}{2})\]

myininaya (myininaya):

\[\cos^{-1}(\frac{-\sqrt{3}}{2})\]

OpenStudy (anonymous):

i forgot the minus sign. sorry. myininaya has it

OpenStudy (anonymous):

yes thats right! now evaluate?

OpenStudy (anonymous):

im supposed to get a degree or something

OpenStudy (anonymous):

ok you sure i dont need a radian or degree?

myininaya (myininaya):

lol ok so do you have your cheat sheet

myininaya (myininaya):

your unit circle

OpenStudy (anonymous):

yes i do

myininaya (myininaya):